Let S={2,3+x,1−x2}, p(x)=2−x−x2 and V=P2

(a) If possible, express p(x)as a linear combination of vectors in S.

(b) By justifying your answer, determine whether the set S is linearly independent or linearly dependent.

(c) By justifying your answer, determine whether the set S is a basis for P2

Please solve it in very detail, and make sure it is correct.
Answer #1

son Q) consider the standcud hasis B = { 1, x,x2} of Pz. Then, the coordinate vectou Weit B are 3 [2]g [3+x3 = 1 [1-x2), = 2we have S= {2, 3+x, 1-2}={v,, V2,3} then consider the malex [vi] B [V₂] B [v3] [bem]8] 2 3 2 0 - 0 -1 O 6 - 3/ /2 RITRA Rz 1-b) from the reduced mates reduced matas we oblained the identity malix to the left. Each collemn of the une mattis correspo